The base edges of a right parallelepiped whose altitude is 8 in., measures 12 in. and

15 in. If the angle included by these edges is 135°, find the volume and the total

surface area of the solid



1
Expert's answer
2021-09-15T02:58:00-0400

Volume "=" L"\times" W"\times" H

8"\times" 12"\times" H"(" cos"\Theta" ")"

8"\times" 12"\times" "(" 0.7071"\times" 15")"

Answer is:

1018.234in"^{3}"

Area;

A"=" 2"(" ab+ac+bc")"

2 "(" 8"\times" 12")" "+(" 8"\times" 15")+" "(" 12"\times" 15")"

2"(" 96 "+" 120"+" 180")"

Answer is;

796in"^{2}"